OpenStudy (anonymous):

Please help! Find the 15th term of the sequence: -3, 6, -12, 24...

OpenStudy (anonymous):

the common ration as it is called is -2

OpenStudy (anonymous):

r=-2

OpenStudy (bahrom7893):

-3 -3*(-2) = 6 6*(-2) = -12 -12*(-2) = 24, etc

OpenStudy (bahrom7893):

Now find the 15th term..

OpenStudy (anonymous):

Got it, thanks guys :)

OpenStudy (anonymous):

-49152

OpenStudy (anonymous):

\[a _{n}=a _{1}(n-1)^{r}\]